OpenStudy (anonymous):

In a game that lasts 48 minutes, exactly 6 players from each team are on the field at all times. Throughout the game, players are substituted so that 8 players on a team each play an equal amount of time. How many minutes is each of the 8 players on the field during the game?

OpenStudy (anonymous):

each player is on the field \[\frac{6}{8}\] of the time, so you want \[\frac{6}{8}\times 48\]